The centre of a circle passing through te points (0,0),(1,0) and touching the circle x^2+y^2=9 is?

Grade:12

4 Answers

shiddhant bhattacharya
11 years ago
i think the
Akash Kumar Dutta
11 years ago

Dear Dipankar,
let m,n be the centre,
then apply deistance formula for the two points mentioned..you get...m=1/2
now we observe that the circle whose centre is required passes through origin.
and also the centre of the other circle is at origin...hence the circles touch each-other internally,
and the radius of the inner circle=1/2.radiu of outer circle=3/2.
now...find n by again using distance formula...you get n=(+-)(2)^1/2
Hence the Centre is [ 1/2 , (+-)root(2) ]
Regards.

sourav singh
11 years ago

[ 1/2, root2 ]

KS
11 years ago

The question seems to be wrong.

As the centre of the given circle is (0,0) , the required circle cannot pass through (0,0) and touch the given circle.